Consider a company with a beta of 1.6. The risk-free interest rate in the market is
4%, whereas the market risk premium is 6%. The recent dividend was just declared for the
company and was $3.00 per share. The required rate of return for the stock is 12%, and the growth
rate of dividends is constant at 5%. A.) Using a dividend discount valuation approach, what is the
intrinsic value of the stock? B.) If you held the stock for one year, and dividends did continue to
grow at 5%, but you also sold the stock after one year for $48.50, instead of its estimated value,
what would be your expected rate of return?
Answer #1

a). According to the CAPM,

Required Return = Risk-free Rate + [Beta*Market Risk Premium]

= 4% + [1.6 * 6%] = 4% + 9.6% = 13.6%

Intrinsic Value = [D0 * (1 + g)] / [r - g]

= [$3 * (1 + 0.05)] / [0.136 - 0.05] = $3.15 / 0.086 = $36.63

b). Required Return = [P1 + Dividends - P0] / P0

= [$48.50 + ($3*1.05) - $36.63] / $36.63

= [$11.87 + $3.15] / $36.63 = $15.02 / $36.63 = 0.4101, or 41.01%